question on permutations and combinations

in two vans, each has 3 seats in the front and 4 at the back. In how many ways 3 girls and 9 boys can be seated in these vans so that the girls sit together at the back.

Grade:

1 Answers

Parvez ali
10 years ago
number of ways of arranging three girls on adjacent back seats of a van= 2*(3!)
thus number of ways of arranging in two vans=2*(2*(3!)=24
on the remaining 11 seats 9 boys has to be arranged by 11p9 ways
thus required number of ways=24*11p9